OpenStudy (anonymous):

Evaluate the integral from sqrt2 to 2 of (1/(t^(3)sqrt(t^(2)-1))dt

OpenStudy (anonymous):

ick. myininaya it is all yours

myininaya (myininaya):

lol make it readable for me please

myininaya (myininaya):

i have to go to the restroom

OpenStudy (anonymous):

ok i will \[\int\frac{1}{t^3\sqrt{t^2-1}}dt\]

OpenStudy (anonymous):

try \[u=\sec(\theta)\] to begin with this is going to be a long one

OpenStudy (anonymous):

\[\int\limits_{\sqrt{2}}^{2} 1\div t ^{3}\sqrt{t ^{2}-1} dx\]

myininaya (myininaya):

ok i'm back

OpenStudy (anonymous):

oh not so bad maybe \[u=\sec(\theta), du=\sec(\theta)\tan(\theta)d\theta \] get \[\int\frac{\sec(\theta)\tan(\theta)d\theta}{\sec^3(\theta)\sqrt{\sec^2(\theta)-1}}\]

myininaya (myininaya):

 let \[\sec(\theta)=t\]